Edelman Engines has $18 billion in total assets — of which cash and equivalents total $110 million. Its balance sheet shows $2.7 billion in current liabilities — of which the notes payable balance totals $1.14 billion. The firm also has $8.1 billion in long-term debt and $7.2 billion in common equity. It has 300 million shares of common stock outstanding, and its stock price is $38 per share. The firm's EBITDA totals $1.482 billion. Assume the firm's debt is priced at par, so the market value of its debt equals its book value. What are Edelman's market/book and its EV/EBITDA ratios? Do not round intermediate calculations. Round your answers to two decimal places.

M/B:

EV/EBITDA:
Answer #1

Book Value (BV) = Total Common Share holders equity / Shares Outstanding    =( 7.2 *10^9) / (300*10^6)

Market / Book Value = stock price / Book Value = 38*(300*10^6)/(7.2*10^9) = 1.58

Enterprise Value ( EV) = Market Capitalization (MC) + Total Debt (TD) - C (Cash and equivalents)

or, EV = MC + Short term debt (Notes Payable)+ Long term Debt -C

= 38*300*10^6 + 1.14*10^9 +8.1*10^9 - 110*10^6

=2.053*10^10

EBITDA = 1.482 * 10 ^9

EV/EBITDA = 2.053*10^10 / 1.482 *10 ^9 = 13.85